OpenStudy (ineedhelpfast12):

Solve the linear system using elimination. 3x-2y=3 -x+y=1

OpenStudy (solomonzelman):

way 1: Multiply the second equation times 3, and add the equations.

OpenStudy (solomonzelman):

way 2: -x+y=1 ---> y=x+1 therefore, 3x-2y=3 ---> 3x-2(x+1)=3 (substitution)

OpenStudy (ineedhelpfast12):

so 3x-2y=3 -3x=3y=3

OpenStudy (solomonzelman):

way 3: multiply the second equation times -2 and add the equations.

OpenStudy (solomonzelman):

, you are going by way 1. ok

OpenStudy (solomonzelman):

check equation 2 please

OpenStudy (ineedhelpfast12):

so y is 1.5?

OpenStudy (solomonzelman):

I don't think so

OpenStudy (solomonzelman):

can you please multiply the second equation times 3, and add the equations.

OpenStudy (ineedhelpfast12):

3x-2y=3 -3x+3y=3

OpenStudy (nincompoop):

first we need to think why are we multiplying

OpenStudy (ineedhelpfast12):

What do you mean?

OpenStudy (nincompoop):

why are you multiplying?

OpenStudy (ineedhelpfast12):

thats the only way for the problem to work. If there is a negative and positive x or y that matches with the other, then the problem works out evenly

OpenStudy (nincompoop):

read the instruction it says solve by elimination. Now what are you trying to eliminate?

OpenStudy (ineedhelpfast12):

either the x or the y

OpenStudy (nincompoop):

good then how are you going to do that?

OpenStudy (ineedhelpfast12):

by multiplying. I dont get the point you are trying to make

OpenStudy (nincompoop):

here is the point, if you knew right from the beginning what you need to do, why are you asking for help?

OpenStudy (ineedhelpfast12):

because I am confused. If you are not going to help me, let me know. My internet shuts off soon, and I need this done by morning

OpenStudy (nincompoop):

I am not saying you cannot ask for help. so you are confused. did you say what part of the solution process you are confused about?

OpenStudy (ineedhelpfast12):

when I multiply and whatever, it cones out to be 2y=3

OpenStudy (ineedhelpfast12):

why is that?

OpenStudy (nincompoop):

okay what were you trying to eliminate?

OpenStudy (ineedhelpfast12):

the x

OpenStudy (nincompoop):

look at the x in the first equation and look at the x in the second equation what are their coefficients?

OpenStudy (ineedhelpfast12):

huh?

OpenStudy (nincompoop):

what is the coefficient of x in the first equation what is the coefficient of x in the second equation 3x-2y=3 (equation1) -x+y=1 (equation 2)

OpenStudy (ineedhelpfast12):

whats a coefficient?

OpenStudy (nincompoop):

get it?

OpenStudy (ineedhelpfast12):

so 3 for equation 1 and -1 for equation 2

OpenStudy (nincompoop):

great

OpenStudy (ineedhelpfast12):

ok?

OpenStudy (nincompoop):